# SVL: Spike-based Vision-language Pretraining for Efficient 3D Open-world Understanding

### v1

@@ 5.3 Ablation Study
An ablation study was conducted to examine the impact of different loss function combinations during our multi-scale triple alignment (MTA). Specifically, we compared performance with and without the semantic spike-text alignment $(e.g.,\mathcal{L}_{(\mathbf{S},\mathbf{T})}^{\text{NCE}})$ and fine-grained spike-image alignment $(e.g.,\mathcal{L}_{(\mathbf{S},\mathbf{I})}^{\text{NCE}}$, $\mathcal{L}_{(\mathbf{S},\mathbf{I})}^{\text{MSE}})$. As shown in Tab. 6, the ablation study highlights the importance of combining loss functions for optimal performance. In the absence of any loss functions, the model only gets 0.5% accuracy on the large-scale Objaverse-LVIS and 5.1% on ModelNet40. Introducing spike-image alignment yields a significant improvement, while the inclusion of semantic spike-text alignment alone demonstrates limited effectiveness. The highest performance is attained when all three loss functions, including the MSE-based fine-grained alignment, are employed, achieving 33.6% accuracy on the large Objaverse-LVIS [41] and 79.6% on ModelNet40 [24]. These findings underscore the synergistic relationship between semantic and fine-grained alignment in enhancing the model’s representational capabilities, showing the effectiveness of our MTA module.
Different Time Steps and Firing Bits Analysis
We systematically evaluated the impact of varying $T$ and $D$ on both SVL pretraining and downstream task fine-tuning, as shown in Tab. 6. For pretraining, we observed that increasing the number of $T$ does not enhance performance, while it adversely affects inference time and significantly increases power consumption. Specifically, after our SVL pretraining, we tested the effect of different $T$ and $D$ on the zero-shot classification accuracy of large-scale Objaverse-LVIS [41]. For instance, with $D=1$, extending $T$ from 2 to 4 resulted in a 0.2% increase in accuracy, accompanied by a doubling of power consumption. Conversely, we found that expanding $D$ while maintaining a fixed $T$ not only improves performance but also reduces power consumption. For fine-tuning, increasing the number of time steps enhances performance but comes at the cost of longer inference time and higher power consumption. In summary, increasing $T$ enhances fine-tuning performance but requires balancing to limit power consumption and inference time, while expanding $D$ offers a solution for improving performance and reducing power demands.
Table 5: Ablation study of the pretrain timesteps.
@@ 6 Conclusion
In this work, we introduce SVL, a spike-based vision-language pretraining framework that equips SNNs with open-world 3D understanding while preserving their inherent energy efficiency. By integrating Multi-scale Triple Alignment (MTA) and a Reparameterizable Vision-Language Integration (Rep-VLI) module, SVL bridges the gap between the low-power advantages of SNNs and the strong generalization capabilities of vision-language models. Comprehensive evaluations across zero-shot 3D classification, semantic segmentation, and human action recognition demonstrate that SVL consistently outperforms prior SNN-based approaches and even rivals state-of-the-art ANNs, all with significantly lower computational cost. Notably, SVL enables SNNs to perform open-world 3D question answering, marking a milestone in multimodal representation learning for spike-based systems. In summary, SVL enables SNNs to acquire powerful, transferable multimodal representations, paving the way for advanced 3D understanding in resource-constrained environments and laying the groundwork for large-scale multimodal learning in neuromorphic computing.
Appendix
Appendix A Backpropagation process of I-LIF
@@
There exist two primary methods of training high-performance SNNs. One way is to discretize ANN into spike form through neuron equivalence [53], i.e., ANN-to-SNN conversion, but this requires a long simulation time step and boosts the energy consumption. We employ the direct training method [29, 54] and apply surrogate gradient training.
@@
Then in this section, we introduce the training process of SNN gradient descent and the parameter update method of spatio-temporal backpropagation (STBP) [29, 55, 56]. SNNs’ parameters can be taught using gradient descent techniques, just like ANNs, after determining the derivative of the generation process. Moreover, the accumulated gradients of loss $\mathcal{L}$ with respect to weights $\mathbf{w}$ at layer $\ell$ can be calculated as:
@@
where ${s}^{\ell}[t]$ and ${u}^{\ell}[t]$ represent the binary and membrane potential of the neuron in layer $\ell$, at time $t$. Moreover, notice that $\frac{\partial{s}^{\ell}[t]}{\partial{u}^{\ell}[t]}$ is non-differentiable. To overcome this problem, [29] propose the surrogate function to make only the neurons whose membrane potentials close to the firing threshold receive nonzero gradients during backpropagation. In this paper, we use the rectangle function, which has been shown to be effective in gradient descent and may be calculated by:
@@
Appendix B Architecture Details
In this section, we present the detailed architectural designs of E-3DSNN, Spike PointNet and Spike PointFormer, outlining their core components, network configurations, and the specific adaptations made to enable efficient analysis of 3D point cloud within the spiking neural network framework.
E-3DSNN [5]
@@ Appendix B Architecture Details
Table 7: Architecture details of E-3DSNN [5]
Spike PointNet [25]
@@
Spike PointFormer
is our proposed Transformer-based SNN backbone for point cloud encoding. Built upon PointBERT [39], we replace ReLU with I-LIF to introduce spiking dynamics. Specifically, the point cloud is first downsampled and grouped into local regions. These are passed through the SNN-adapted encoder to extract local features. Then, Spike PointFormer stacks $L$ spike-driven Transformer layers, enabling global feature interaction. Together, Spike PointFormer can be formulated as:
$x={\rm{KNN}}(\rm{FPS}(P)),$ $P\in\mathbb{R}^{B\times N\times 3},\quad x\in\mathbb{R}^{B\times N^{\prime}\times K\times 3}$ (13)
$f_{0}={\rm{EMP}}({\rm{MLP}}({\mathcal{SN}}\left(x\right))),$ $f_{0}\in\mathbb{R}^{B\times N^{\prime}\times D}$ (14)
$f_{l}={\rm{SDT}}(f_{l-1})+f_{l-1},$ $f_{l}\in\mathbb{R}^{B\times N^{\prime}\times D},l=1...L$ (15)
where $P\in\mathbb{R}^{B\times N\times 3}$ is the input point cloud with batch size $B$, point number $N$. ${\rm FPS}(\cdot)$ and ${\rm KNN}(\cdot)$ denote Farthest Point Sampling and K-nearest Neighbor Grouping. ${\mathcal{SN}}(\cdot)$ represents spike neuron; ${\rm MLP}(\cdot)$ consists of multiple convs; ${\rm EMP}(\cdot)$ denotes Element-wise Max Pooling. ${\rm SDT}(\cdot)$ denotes Spike-Driven Transformer layer. $f_{0}$ and $f_{l}$ are intermediate features with dimension $D$.
Notably, to better align with SNN principles, we remove the softmax in the spike-driven transformer layer and re-arrange the computation order of Query, Key and Value to achieve sparse and energy-efficient attention. In particular, the spike-driven attention mechanism can be formulated as:
$\operatorname{SDA}\left(Q,K,V\right)=\mathcal{SN}\left(\mathcal{SN}\left(Q\right)\odot\mathcal{SN}\left(K\right)^{\mathrm{T}}\right)\odot\mathcal{SN}\left(V\right),$ (16)
where $\operatorname{SDA}(\cdot)$ denotes the spike-driven attention, and $\odot$ represents the dot-product operation. $\mathcal{SN}(\cdot)$ represents spike neurons referred to I-LIF.
Appendix C Open-world Multimodel Learning Details
Figure 3: Architecture details of open-world multimodel learning.
In this section, we present the details of how to perform open-world multimodal learning after pretraining with the SVL model. Our primary goal is to effectively leverage the capabilities of the pretrained LLM and the spike-based encoder pretrained with SVL. The network architecture is shown in Fig. 3. We select vicuna [44] as our LLM $\mathcal{F}_{\theta}(\cdot)$, parameterized by $\theta$, use the Spike Pointformer $\mathcal{E}^{S}_{\theta}(\cdot)$ pretrained with SVL as the spiking visual encoder, and the projector $\mathcal{P}_{\theta}(\cdot)$.
$V^{t}=\mathcal{P}_{\theta}(\mathcal{E}^{S}_{\theta}(D^{t}_{k})),$ (17)
$R^{t}=\mathcal{F}_{\theta}(I^{t};V^{t}),$ (18)
Table 8: Qualitative comparisons. We show the qualitative results of models on the ScanNet [57]. Our SVL-13B can understand 3D semantics and respond to prompts effectively comparable to other ANN-based models.
Table 9: Qualitative comparisons. We show the qualitative results of models on ModelNet40 [24] and Objaverse [41]. Our SVL-13B can understand 3D semantics and respond to prompts effectively comparable to other ANN-based models.
       Ground Truth         Laptop         A cartoon black monster like a dragon
       Prompt         What is this?         Briefly caption this 3D model.
       InstructBLIP [43]         symbol letter l         a black lizard with a sharp tooth in a dark room
       LLaVA [45]         A small, grainy, black and white letter j.         A 3D model of a dark, menacing dragon.
       3D-LLM [58]         -         A black and white tiger with long legs, standing on its hind leg.
       Point-Bind LLM [59]         This is a laptop computer.         The 3D model features a large, ornate gargoyle with a horned helmet, sitting on top of a building.
       PointLLM [12]         The 3D model represents a notebook computer, typically a laptop.         The 3D model depicts a menacing black dragon, with its mouth opened wide revealing a row of sharp teeth.
Appendix D Theoretical Energy Consumption
In our SVL framework, Rep-VLI can transform the text embeddings into tiny weights during inference. Additionally, the framework can convert matrix multiplication into sparse addition, which can be implemented as addressable additions on neuromorphic chips. In the first coding layer, convolution operations act as Multiply-Accumulate (MAC) operations that convert analog inputs into spikes, similar to direct coding-based SNNs [60]. Similarly, in the final layer, logit calculations also perform MAC operations. In contrast, in the SNN architecture, the convolution (Conv) or fully connected (FC) layers transmit spikes and execute Accumulation (AC) operations to accumulate weights for postsynaptic neurons. Hence, the inference energy cost for our SVL framework can be expressed as follows:
where $N$ and $M$ represent the total number of sparse spike convolutions, and $E_{MAC}$ and $E_{AC}$ are the energy costs associated with MAC and AC operations, respectively. The variables $fr^{m}$, $fr^{n}$, $FL_{conv}^{n}$, and $FL_{fc}^{m}$ denote the firing rate and FLOPs of the $n$-th sparse spike convolution layer. Previous SNN studies [61, 62, 54, 63] assume a 32-bit floating-point implementation in 45nm technology, with $E_{MAC}$ = 4.6 pJ and $E_{AC}$ = 0.9 pJ for various operations.
Appendix E Implementation Details
Table 10: Hyper-parameters for SVL pretraining.
Hyper-parameter E-3DSNN-T/S/L/H Spike PointNet Spike PointFormer
Timestep (Training/Inference) $1\times 4$/ $4\times 1$ $1\times 4$/ $4\times 1$ $1\times 4$/ $4\times 1$
Epochs 250 250 250
Batch size 4096 1024 1024
Optimizer AdamW AdamW AdamW
Base Learning rate $32e-$ 3 $32e-$ 3 $33e-$ 3
Learning rate decay Cosine Cosine Cosine
Warmup eopchs 10 10 10
Weight decay $41e-$ 4 $41e-$ 4 0.1
Table 11: Hyper-parameters for SVL Finetuning on 3D point cloud.
Hyper-parameter ModelNet40 ScanObjectNN KITTI SemanticKITTI
Timestep (Training/Inference) $1\times 4$/ $4\times 1$ $1\times 4$/ $4\times 1$ $1\times 4$/ $4\times 1$ $1\times 4$/ $4\times 1$
Epochs 300 250 100 80
Batch size 64 64 96 64
Optimizer AdamW AdamW AdamW AdamW
Base Learning rate $45e-$ 4 $35e-$ 3 $21e-$ 2 $32e-$ 3
Learning rate decay Onecycle Onecycle Onecycle Onecycle
Table 12: Hyper-parameters for SVL Finetuning on DVS datasets.
Hyper-parameter DVS Action DVS128 Gesture
Timestep (Training/Inference) $1\times 4$/ $6\times 4$ $1\times 4$/ $6\times 4$
Epochs 250 250
Batch size 4096 1024
Optimizer AdamW AdamW
Base Learning rate $32e-$ 3 $32e-$ 3
Learning rate decay Cosine Cosine
Warmup eopchs 10 10
Weight decay $41e-$ 4 $41e-$ 4
Appendix F Datasets
[…]
The Objaverse dataset, which includes Objaverse-LVIS [41] as a subset, is currently the largest 3D dataset. Objaverse-LVIS is a significant part of the Objaverse dataset, containing 46,832 annotated shapes across 1,156 LVIS categories. This extensive collection of 3D shapes provides a rich resource for researchers and practitioners in the field of computer vision and 3D modeling.
The large KITTI dataset [64] contains 7481 training samples, 3717 of which constitute trainsets and 3769 of which constitute validation sets. E-3DSNN is evaluated as backbones equipped with VoxelRCNN Head In detection [65]. To execute our model, we uses OpenPCDet [66] that is transformed into a spiking version by us. After being divided into regular voxels, raw point clouds are input to our 3DSNN on KITTI [48]. The point clouds are clipped to ranges of $[-0.7\text{m},0.4\text{m}]$ for the X-axis, $[-40\text{m},40\text{m}]$ for the Y-axis, and $[-3\text{m},1\text{m}]$ for the Z-axis followed by voxelization with a resolution of $(0.05\text{m},0.05\text{m},0.1\text{m})$. The Average Precision (AP) calculated by 11 recall positions for the Car class is used as the evaluation metrics.
[…]
The DVS Action dataset [50] comprises 10 actions performed by 15 subjects within 5s, which is recorded by DVS camera in an empty office. DVS is a vision sensor [50] that can records a sequence of tuples $[t,x,y,p]$ for each event streams. Among them, $t$ represents the timestamp of the event, $(x,y)$ represents the event’s pixel coordinates and $p$ represents the polartity of the event.
The DVS128 Gesture dataset [49] contains 1,342 instances across 11 different hand and arm gestures, which are performed by 29 subjects under 3 distinct lighting conditions in 122 trials. They are captured by DVS128 camera, a DVS with $128\times 128$ pixel resolution.
Appendix G Limitations
One of the limitations of our work is exploring scalability of E-3DSNN. The largest model reported is E-3DSNN-H with 46.7M parameters. Future work will explore the performance and computational efficiency trade-offs of SVL at larger parameter scales.
The experimental results presented in this paper are reproducible. Detailed explanations of model training and configuration are provided in the main text and supplemented in the appendix. Our codes and models will be made available on GitHub after review.

### v2

@@ 5.3 Ablation Study
An ablation study was conducted to examine the impact of different loss function combinations during our multi-scale triple alignment (MTA). Specifically, we compared performance with and without the semantic spike-text alignment $(e.g.,\mathcal{L}_{(\mathbf{S},\mathbf{T})}^{\text{NCE}})$ and fine-grained spike-image alignment $(e.g.,\mathcal{L}_{(\mathbf{S},\mathbf{I})}^{\text{NCE}}$, $\mathcal{L}_{(\mathbf{S},\mathbf{I})}^{\text{MSE}})$. As shown in Tab. 6, the ablation study highlights the importance of combining loss functions for optimal performance. In the absence of any loss functions, the model only gets 0.5% accuracy on the large-scale Objaverse-LVIS and 5.1% on ModelNet40. Introducing spike-image alignment yields a significant improvement, while the inclusion of semantic spike-text alignment alone demonstrates limited effectiveness. The highest performance is attained when all three loss functions, including the MSE-based fine-grained alignment, are employed, achieving 33.6% accuracy on the large Objaverse-LVIS (Deitke et al., 2023) and 79.6% on ModelNet40 (Wu et al., 2015). These findings underscore the synergistic relationship between semantic and fine-grained alignment in enhancing the model’s representational capabilities, showing the effectiveness of our MTA module.
@@
Different Time Steps and Firing Bits
We systematically study time steps ($T$) and firing bits ($D$) for SVL pretraining and downstream fine-tuning (Tab. 7). During pretraining, enlarging $T$ offers negligible accuracy gains yet degrades efficiency: for example, with D=1D{=}1, increasing $T$ from 2 to 4 improves Objaverse-LVIS zero-shot accuracy by only 0.2% while roughly doubling power and worsening latency. In contrast, scaling $D$ at a fixed $T$ consistently improves accuracy and can even reduce power by concentrating information into fewer, stronger spikes. For fine-tuning, higher $T$ can help recognition quality, but the benefit comes with longer inference and higher energy. In practice, we recommend small $T$ (often T=1T{=}1) with moderate $D$ for pretraining to minimize compute, then using modest $T$ (e.g., 2–4) and tuning $D$ for downstream tasks to balance accuracy against latency and power.
Table 7: Ablation study of the pretrain timesteps.
@@ 6 Conclusion
In this work, we introduce SVL, a novel spike-based vision–language pretraining framework designed to equip Spiking Neural Networks with robust open-world 3D understanding capabilities while simultaneously preserving their inherent energy efficiency. Through the strategic integration of Multi-scale Triple Alignment and a Reparameterizable Vision–Language Integration module, our approach effectively bridges the longstanding gap between the low-power advantages characteristic of SNNs and the strong generalization capabilities typically associated with advanced vision–language models. Our comprehensive evaluations, which span diverse tasks including zero-shot 3D classification, semantic segmentation, and human action recognition, demonstrate that SVL consistently outperforms prior SNN-based approaches. Furthermore, it achieves performance levels that rival state-of-the-art Artificial Neural Networks, all while operating at a significantly lower computational cost. Notably, SVL empowers SNNs with the unprecedented ability to perform open-world 3D question answering, which marks a significant milestone in the field of multimodal representation learning for spike-based systems. Finally, we present the Spike-driven PointFormer, which stands as the first fully spike-driven point Transformer. This architecture features a 3D spike-driven self-attention mechanism that reduces complex interactions to sparse additions, thereby delivering faster and more energy-efficient training dynamics while maintaining strong predictive accuracy.
Acknowledgements
This work was partially supported by the Zhongguancun Academy (Grant No.s 1800302002), National Distinguished Young Scholars (62325603), National Natural Science Foundation of China (62236009, U22A20103, U2541222), CAS Project for Young Scientists in Basic Research (YSBR116), Beijing Science and 698 Technology Plan (Z241100004224011), and CAAITencent Rhino-Bird Open Research Fund.
Impact Statement
This paper advances the development of SNNs for open–world 3D understanding, bridging the gap between SNNs and ANNs while preserving energy efficiency. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
Appendix A More Experiments
This section augments the main results with three complementary studies: comparisons between different Backbone, dynamic scene segmentation on Synthia 4D (Ros et al., 2016) to assess temporal generalization, ablations on event/discretization design (voxel size and point density), and the impact of CLIP image‐encoder size used to build vision–language prototypes. For each study, we report results and provide a concise analysis.
Figure 3: Overall architecture of our Spike-driven PointFormer.
A.1 Speedup across spike point Transformers
[…]
Table 8: Ablation study of different backbone efficiency on ModelNet40.
[…]
A.2 Dynamic Scene Segmentation on Synthia 4D
We evaluate temporal scene understanding on Synthia 4D using voxelized inputs and multi‐step spike simulation. SVL improves the SNN backbone and matches or surpasses ANN counterparts under comparable capacity.
Table 9: Scene segmentation on Synthia 4D (Ros et al., 2016). SVL enables strong temporal understanding for SNNs.
Method Input Frames Params (M) mIoU (%)
3D MinkNet14 (Ros et al., 2016) Voxel 1 19.3 76.24
4D MinkNet14 (Ros et al., 2016) Voxel 3 23.7 77.46
Same-structure ANN Voxel 3 19.1 79.54
E-3DSNN-L (w/o SVL) Voxel $3\times 2$ 19.1 76.41
E-3DSNN-L (SVL, ours) Voxel $1\times 2$ 17.7 78.91
E-3DSNN-L (SVL, ours) Voxel $3\times 2$ 19.1 80.05
SVL yields a substantial gain over the SNN backbone without pretraining (80.05 vs. 76.41 mIoU) and slightly surpasses a capacity‐matched ANN (79.54 mIoU), indicating that aligning spike features to the vision–language space improves scene‐level generalization; moreover, increasing temporal extent from $1\times 2$ to $3\times 2$ frames further boosts performance (78.91 to 80.05 mIoU), suggesting that spike‐driven temporal accumulation is effectively exploited.
A.3 Event Construction Ablations
We ablate voxel size and the number of input points on ModelNet40 zero‐shot with E-3DSNN-T + SVL.
Table 10: Ablations on discretization. Moderate voxel size ($0.02$) and sufficient point density (10k–20k) work best.
Voxel size 0.01 0.02 0.04
Top-1 Acc. (%) 78.9 79.6 79.1
#Points 5k 10k 20k
Top-1 Acc. (%) 77.3 79.6 79.8
Accuracy peaks at a moderate voxel size of 0.02, where overly fine discretization (0.01) fragments geometry and increases sparsity noise while overly coarse discretization (0.04) washes out structure; likewise, raising point density from 5k to 10k yields a clear improvement whereas the gain from 10k to 20k is marginal, indicating diminishing returns and suggesting that a mid‐range density strikes the best balance between fidelity and efficiency.
A.4 CLIP Encoder Size
We study the impact of the image encoder size used to build vision–language prototypes during pretraining.
Table 11: Larger vision encoders yield stronger zero-shot 3D alignment.
Image encoder ModelNet40 Top-1 (%) Objaverse-LVIS Top-1 (%)
OpenCLIP ViT-B 66.8 25.1
OpenCLIP ViT-G 70.6 27.9
OpenCLIP ViT-bigG 79.6 33.6
Scaling the vision encoder substantially strengthens zero‐shot transfer (e.g., ViT-B to bigG: +12.8 on ModelNet40 and +8.5 on Objaverse-LVIS), reflecting richer visual semantics that better anchor the spike encoder in the shared space; notably, because Rep‐VLI removes the text encoder at inference, larger CLIP backbones increase precomputation and training cost but do not affect runtime latency or energy, making bigG preferable when resources allow.
Appendix B Open-world Multimodel Learning Details
Figure 4: Architecture details of open-world multimodel learning.
In this section, we present the details of how to perform open-world multimodal learning after pretraining with the SVL model. Our primary goal is to effectively leverage the capabilities of the pretrained LLM and the spike-based encoder pretrained with SVL. The network architecture is shown in Fig. 4. We select vicuna (Chiang et al., 2023) as our LLM $\mathcal{F}_{\theta}(\cdot)$, parameterized by $\theta$, use the Spike-driven Pointformer $\mathcal{E}^{S}_{\theta}(\cdot)$ pretrained with SVL as the spiking visual encoder, and the projector $\mathcal{P}_{\theta}(\cdot)$.
$V^{t}=\mathcal{P}_{\theta}(\mathcal{E}^{S}_{\theta}(D^{t}_{k})),$ (17)
$R^{t}=\mathcal{F}_{\theta}(I^{t};V^{t}),$ (18)
Table 12: Qualitative comparisons. We show the qualitative results of models on the ScanNet (Dai et al., 2017). Our SVL-13B can understand 3D semantics and respond to prompts effectively comparable to other ANN-based models.
Table 13: Qualitative comparisons. We show the qualitative results of models on ModelNet40 (Wu et al., 2015) and Objaverse (Deitke et al., 2023). Our SVL-13B can understand 3D semantics and respond to prompts effectively comparable to other ANN-based models.
   Ground Truth     Laptop     A cartoon black monster like a dragon
   Prompt     What is this?     Briefly caption this 3D model.
   InstructBLIP (Dai et al., 2023)     symbol letter l     a black lizard with a sharp tooth in a dark room
   LLaVA (Liu et al., 2023a)     A small, grainy, black and white letter j.     A 3D model of a dark, menacing dragon.
   3D-LLM (Hong et al., 2023)     -     A black and white tiger with long legs, standing on its hind leg.
   Point-Bind LLM (Guo et al., 2023)     This is a laptop computer.     The 3D model features a large, ornate gargoyle with a horned helmet, sitting on top of a building.
   PointLLM (Xu et al., 2024)     The 3D model represents a notebook computer, typically a laptop.     The 3D model depicts a menacing black dragon, with its mouth opened wide revealing a row of sharp teeth.
Appendix C Theoretical Energy Consumption
In our SVL framework, Rep-VLI can transform the text embeddings into tiny weights during inference. Additionally, the framework can convert matrix multiplication into sparse addition, which can be implemented as addressable additions on neuromorphic chips. In the first coding layer, convolution operations act as Multiply-Accumulate (MAC) operations that convert analog inputs into spikes, similar to direct coding-based SNNs (Wu et al., 2019). Similarly, in the final layer, logit calculations also perform MAC operations. In contrast, in the SNN architecture, the convolution (Conv) or fully connected (FC) layers transmit spikes and execute Accumulation (AC) operations to accumulate weights for postsynaptic neurons. Hence, the inference energy cost for our SVL framework can be expressed as follows:
where $N$ and $M$ represent the total number of sparse spike convolutions, and $E_{MAC}$ and $E_{AC}$ are the energy costs associated with MAC and AC operations, respectively. The variables $fr^{m}$, $fr^{n}$, $FL_{conv}^{n}$, and $FL_{fc}^{m}$ denote the firing rate and FLOPs of the $n$-th sparse spike convolution layer. Previous SNN studies (Horowitz, 2014; Rathi and Roy, 2021; Qiu et al., 2024, 2023) assume a 32-bit floating-point implementation in 45nm technology, with $E_{MAC}$ = 4.6 pJ and $E_{AC}$ = 0.9 pJ for various operations.
Appendix D Implementation Details
Table 14: Hyperparameters for SVL pretraining.
Architecture E-3DSNN-T/S/L/H Spike PointNet Spike-driven PointFormer Timestep (Training/Inference) $1\times 4$/ $4\times 1$ $1\times 4$/ $4\times 1$ $1\times 4$/ $4\times 1$ Epochs 250 250 250 Batch size 4096 1024 1024 Optimizer AdamW AdamW AdamW Base Learning rate 2e32e-3 2e32e-3 3e33e-3 Learning rate decay Cosine Cosine Cosine Warmup eopchs 10 10 10 Weight decay 1e41e-4 1e41e-4 0.1
Table 15: Hyper-parameters for SVL Finetuning on 3D point cloud.
Hyper-parameter ModelNet40 ScanObjectNN KITTI SemanticKITTI Timestep (Training/Inference) $1\times 4$/ $4\times 1$ $1\times 4$/ $4\times 1$ $1\times 4$/ $4\times 1$ $1\times 4$/ $4\times 1$ Epochs 300 250 100 80 Batch size 64 64 96 64 Optimizer AdamW AdamW AdamW AdamW Base Learning rate 5e45e-4 5e35e-3 1e21e-2 2e32e-3 Learning rate decay Onecycle Onecycle Onecycle Onecycle
Table 16: Hyper-parameters for SVL Finetuning on temporal datasets.
Hyper-parameter Synthia 4D DVS Action DVS128 Gesture
Timestep (Training/Inference) $1\times 2$/ $3\times 2$ $1\times 4$/ $6\times 4$ $1\times 4$/ $6\times 4$
Epochs 250 250 250
Batch size 64 64 64
Optimizer AdamW AdamW AdamW
Base Learning rate 2e32e-3 2e32e-3 2e32e-3
Learning rate decay Cosine Cosine Cosine
Weight decay 1e41e-4 1e41e-4 1e41e-4
Table 17: Hyper-parameters and training details for Spike-driven PointFormer with SVL on open-world multimodal learning.
Hyper-parameter Stage-1 (Feature Alignment) Stage-2 (Instruction Tuning)
Optimizer AdamW AdamW
Learning rate decay Cosine Cosine
Epochs 3 3
Batch size 128 32
Base Learning rate $2\times 10^{-3}$ $2\times 10^{-5}$
Weight decay $\times 10^{-4} 1×1$$\times 10^{-4}$
Dataset 660K 70K
Appendix E Limitations
This study also has several limitations, which simultaneously highlight promising avenues for further work. First, while SVL demonstrates substantial efficiency gains, our current evaluation is restricted to specific model scales. Validating the scalability of the proposed spike-driven mechanisms on larger-scale foundation models remains a crucial step to fully establish their applicability to more complex tasks.
Appendix F Datasets
The ModelNet40 (Wu et al., 2015) dataset contains 12,311 CAD models across 40 object categories. Among them, 9,843 models are for training and 2,468 are for testing. The point clouds are clipped to ranges of $0.2\text{m},0.2\text{m}]$ for all X-, Y-, and Z-axes as the input data followed by voxelization with a resolution of $0.01\text{m}$. Classification performance was measured using overall accuracy metrics.
ScanObjectNN (Uy et al., 2019) consists of 11,416 training and 2,882 testing samples of real-world scanned 3D objects across 15 categories, with different degrees of data missing and noise contamination. The point clouds are clipped to ranges of $0.2\text{m},0.2\text{m}]$ for all X-, Y-, and Z-axes as the input data followed by voxelization with a resolution of $0.01\text{m}$.
The Objaverse dataset, which includes Objaverse-LVIS (Deitke et al., 2023) as a subset, is currently the largest 3D dataset. Objaverse-LVIS is a significant part of the Objaverse dataset, containing 46,832 annotated shapes across 1,156 LVIS categories. This extensive collection of 3D shapes provides a rich resource for researchers and practitioners in the field of computer vision and 3D modeling.
The large KITTI dataset (Geiger et al., 2012a) contains 7481 training samples, 3717 of which constitute trainsets and 3769 of which constitute validation sets. E-3DSNN is evaluated as backbones equipped with VoxelRCNN Head In detection (Deng et al., 2021). To execute our model, we uses OpenPCDet (Team, 2020) that is transformed into a spiking version by us. After being divided into regular voxels, raw point clouds are input to our 3DSNN on KITTI (Geiger et al., 2012b). The point clouds are clipped to ranges of $0.7\text{m},0.4\text{m}]$ for the X-axis, $40\text{m},40\text{m}]$ for the Y-axis, and $3\text{m},1\text{m}]$ for the Z-axis followed by voxelization with a resolution of $(0.05\text{m},0.05\text{m},0.1\text{m})$. The Average Precision (AP) calculated by 11 recall positions for the Car class is used as the evaluation metrics.
The large SemanticKITTI dataset (Behley et al., 2019) contains 22 sequences from the raw KITTI dataset. About 1,000 lidar scans are included in each sequence, each of which corresponds to approximately 20,000 individual frames. We first adapted the Pointcept (Contributors, 2023) codebase into a spiking neural network (SNN) framework and utilized this customized implementation for model execution. Subsequently, we designed an asymmetric encoder-decoder architecture inspired by the UNet (Choy et al., 2019; Wu et al., 2023) paradigm, where the E-3DSNN acts as the encoder to extract hierarchical multi-scale features, while the decoder progressively fuses these features through skip connections to refine the output. During voxelize implementation, we set the window size to $120\text{m},2^{\circ},2^{\circ}]$ for $(r,\theta,\phi)$. For data preprocessing, the input scene is restricted to the range $51.2\text{m},-51.2\text{m},-4\text{m}]$ to $51.2\text{m},51.2\text{m},2.4\text{m}]$. The voxel size is set to $0.1\text{m}$.
The DVS Action dataset (Miao et al., 2019) comprises 10 actions performed by 15 subjects within 5s, which is recorded by DVS camera in an empty office. DVS is a vision sensor (Miao et al., 2019) that can records a sequence of tuples $[t,x,y,p]$ for each event streams. Among them, $t$ represents the timestamp of the event, $(x,y)$ represents the event’s pixel coordinates and $p$ represents the polartity of the event.
The DVS128 Gesture dataset (Amir et al., 2017) contains 1,342 instances across 11 different hand and arm gestures, which are performed by 29 subjects under 3 distinct lighting conditions in 122 trials. They are captured by DVS128 camera, a DVS with $128\times 128$ pixel resolution.
Synthia 4D. We employ the Synthia dataset (Ros et al., 2016) to construct 3D video sequences. Specifically, we use six driving scenarios across nine different weather conditions. Each scenario provides four stereo RGB-D images captured from the roof of a moving car. Depth images are back-projected into 3D space to generate 3D video sequences. For training, we use sequences 1–4, excluding the sunset, spring, and fog conditions; validation is conducted on sequence 5 under foggy weather; and testing is performed on sequence 6 under sunset and spring conditions. In total, the training, validation, and test splits contain 20,000, 815, and 1,886 3D scenes, respectively. Since the dataset is fully synthetic, we augment it with various types of noise to simulate realistic observations. These include elastic distortion, Gaussian noise, and chromatic shifts applied to the input point clouds.
MLLM training dataset. We further construct a large-scale point–text instruction-following dataset comprising approximately 730K samples and 60K instruct dataset following (Xu et al., 2024). This dataset is designed to support effective training by covering a broad spectrum of topics such as color, shape, usage, and material, thereby enabling robust multimodal instruction-following capabilities.
Appendix G Backpropagation process of I-LIF
@@
There exist two primary methods of training high-performance SNNs. One way is to discretize ANN into spike form through neuron equivalence (Li et al., 2021b), i.e., ANN-to-SNN conversion, but this requires a long simulation time step and boosts the energy consumption. We employ the direct training method (Wu et al., 2018; Qiu et al., 2024) and apply surrogate gradient training.
@@
Then in this section, we introduce the training process of SNN gradient descent and the parameter update method of spatio-temporal backpropagation (STBP) (Wu et al., 2018; Xiao et al., 2022; Hu et al., 2024). SNNs’ parameters can be taught using gradient descent techniques, just like ANNs, after determining the derivative of the generation process. Moreover, the accumulated gradients of loss $\mathcal{L}$ with respect to weights $\mathbf{w}$ at layer $\ell$ can be calculated as:
@@
where ${s}^{\ell}[t]$ and ${u}^{\ell}[t]$ represent the binary and membrane potential of the neuron in layer $\ell$, at time $t$. Moreover, notice that $\frac{\partial{s}^{\ell}[t]}{\partial{u}^{\ell}[t]}$ is non-differentiable. To overcome this problem, (Wu et al., 2018) propose the surrogate function to make only the neurons whose membrane potentials close to the firing threshold receive nonzero gradients during backpropagation. In this paper, we use the rectangle function, which has been shown to be effective in gradient descent and may be calculated by:
@@
Appendix H Architecture Details
In this section, we present the detailed architectural designs of E-3DSNN, Spike PointNet and Spike-driven PointFormer, outlining their core components, network configurations, and the specific adaptations made to enable efficient analysis of 3D point cloud within the spiking neural network framework.
E-3DSNN (Qiu et al., 2025a)
@@ Appendix H Architecture Details
Spike PointNet (Ren et al., 2024)
@@
Spike-driven PointFormer
is our proposed Transformer-based SNN backbone for point cloud encoding.